# pFedMMA: Personalized Federated Fine-Tuning with Multi-Modal Adapter for Vision-Language Models

### v1

@@ 5.2 Performance Evaluation
Base-to-Novel Class Generalization. We assess the performance of pFedMMA in terms of its ability to generalize from local training classes to both base and novel classes. Following prior work, we report the top-1 accuracy on each client’s local classes, on the base classes seen by other clients, and on novel classes that are entirely unseen during training. To capture the overall effectiveness, we use the harmonic mean (HM) of these three metrics. As summarized in Table 1, pFedMMA consistently achieves strong performance across all evaluation categories. It delivers the best overall harmonic mean across the average of seven datasets, outperforming all baselines. Notably, pFedMMA achieves the highest novel class accuracy and matches the top result for base class accuracy, demonstrating excellent generalization. On local classes, it remains competitive, only slightly behind FedPGP, FedOTP, and pFedMoAP.
Compared to PromptFL, which tends to favor generalization at the expense of personalization, pFedMMA achieves a more balanced performance. FedOTP shows high local accuracy in some cases, but its extremely low base and novel class scores indicate poor generalization. pFedMoAP performs well on local classes due to its MoE-based prompt sharing, but lags behind pFedMMA on novel and base class accuracy. These results highlight pFedMMA’s strength in balancing personalization with strong generalization, particularly in scenarios with non-overlapping class distributions across clients.
Model Evaluation on Feature & Label Shifts. To assess the robustness of pFedMMA under realistic federated learning scenarios, we evaluate its performance under both label shift and feature shift using the DomainNet and Office-Caltech10 datasets. In this setting, each domain is divided into two clients using a Dirichlet distribution with $\beta=0.3$, resulting in a total of $8$ clients for DomainNet and $12$ clients for Office-Caltech10. The results are summarized in Table 2. Under these challenging heterogeneous settings, traditional methods such as PromptFL and CLIP show limited ability to generalize, with PromptFL achieving only $12.63$ percent average accuracy across Office-Caltech10 and DomainNet. In contrast, pFedMMA outperforms all baselines, achieving the highest average accuracy of $31.49$ percent. These results highlight the effectiveness and generalization capability of pFedMMA in scenarios with both label and feature heterogeneity, which are common in real-world federated applications.
@@
Evaluation on Personalization. We further evaluate the personalization capability of pFedMMA on the CIFAR-10 and CIFAR-100 datasets under a challenging Dirichlet partitioning scheme with $\beta=0.3$ across $100$ clients and $10$ percent client participation per communication round. The results are presented in Table 3. Compared to the baseline methods, pFedMMA achieves consistently higher accuracy on both datasets, highlighting its strong ability to adapt to highly non-IID data distributions.
Learning Curves. To further investigate the convergence behavior of pFedMMA, we visualize the test accuracy over communication rounds across four representative datasets in Fig 2. Each method is evaluated under the same federated setup with local training epoch $2$ and communication round $50$. As shown, pFedMMA achieves consistently high accuracy and demonstrates stable and fast convergence across all datasets. Notably, despite pFedMMA achieving superior generalization on both base and novel classes (as shown in Table 1), its local performance is still comparable to pFedMoAP throughout the training process. These results confirm that pFedMMA maintains strong personalization while steadily improving generalization.
Figure 2: Accuracy learning curves of pFedMMA and baselines on four datasets over $10$ clients.

### v2

@@ 4.2 Performance Evaluation
Base-to-Novel Class Generalization. We evaluate the performance of pFedMMA in terms of its ability to generalize from locally trained classes to both base and novel classes. Following prior work, we report top-1 accuracy on each client’s local classes, on the base classes seen by other clients, and on novel classes that are entirely unseen during training. To capture overall effectiveness, we use the harmonic mean (HM) of these three metrics. As summarized in Table 1 for the 16-shot setting, pFedMMA consistently achieves strong performance across all evaluation categories and delivers the best overall HM averaged across seven datasets, outperforming all baselines.
Zero-shot CLIP and PromptFL suffer from poor local accuracy, tending to favor generalization at the expense of personalization. We also report $\Delta$, which denotes the relative improvement of pFedMMA compared with the strongest non-baseline methods (FedPGP, FedOTP, and pFedMoAP). While FedOTP sometimes achieves high local accuracy, its extremely low base and novel class scores indicate poor generalization. pFedMoAP performs well on local classes due to its MoE-based prompt sharing, but it lags behind pFedMMA in base and novel accuracy. By contrast, pFedMMA achieves the highest base and novel accuracy, surpassing FedPGP and demonstrating excellent generalization, while remaining competitive on local classes—only $0.74\%$ lower than pFedMoAP.
Fig. 3 illustrates local and HM accuracy across varying numbers of shots $\{1,2,4,8,16\}$, showing the same performance pattern. Detailed results for all datasets are provided in Table 7 in the Appendix.
@@
Evaluation on Personalization. We further evaluate the personalization capability of pFedMMA on CIFAR-10 and CIFAR-100 under a challenging Dirichlet partitioning scheme, varying the concentration parameter $\beta$ across 100 clients with 10% client participation per communication round. The results, summarized in Table 2, show that pFedMMA consistently achieves the highest accuracy on both datasets, demonstrating its strong adaptability to highly non-IID data distributions.
Model Evaluation on Feature & Label Shifts. To evaluate the robustness of pFedMMA in realistic federated learning scenarios, we examine its performance under both label shift and feature shift using the DomainNet and Office-Caltech10 datasets. Following the standard protocol, each domain is split into two clients via a Dirichlet distribution with $\beta=0.5$, yielding 12 clients for DomainNet and 8 clients for Office-Caltech10. The results in Table 3 show that under these challenging heterogeneous conditions, traditional methods such as CLIP and PromptFL struggle to generalize effectively. In contrast, pFedMMA consistently achieves the highest average accuracy across both datasets, highlighting its strong robustness to cross-domain shifts. Additional experiments with one or two clients per domain and varying $\beta$ are provided in Tables 9 and 10 in the Appendix.
4.3 Ablation Study
Impact of model. To further examine the performance of pFedMMA under a different backbone, we report results with ViT-B/32 on the average of six datasets across five shot settings, comparing against three advanced baselines (Table 4). While pFedMMA shows slightly lower local accuracy than FedOTP and pFedMoAP, this gap narrows as the number of shots increases. Importantly, pFedMMA consistently achieves the best trade-off between personalization and generalization, demonstrating stable improvements in the harmonic mean across all settings. Detailed results for all datasets are provided in Table 8 in the Appendix.
Dimension of the Shared Layer. Table 5 (bottom-left) reports the average accuracies over four datasets and five shot settings. As shown, using a larger 128-dimensional representation yields slightly better performance than 32 dimensions. However, to keep the number of trainable parameters low, we consistently adopt the 32-dimensional setting throughout the paper. Detailed results are provided in Tables 11 and 12 in the Appendix.
Scaling Factor $\alpha$. The scaling factor controls the balance between general features and task-specific features. We systematically evaluate its effect, with results summarized in Table 5 (top-left). Our pFedMMA achieves the best trade-off performance (HM) between local, base, and novel classes at $\alpha=0.005$. A larger scaling factor enables faster adaptation to base classes but leads to weaker performance on novel and base classes, whereas a smaller scaling factor hinders effective tuning for downstream tasks. Detailed results are provided in Table 14 in the Appendix.
Starting Layer $\ell$. We evaluate different choices of encoder layers for integrating pFedMMA in Table 5 (top-right). As shown, updating the last three layers yields the best HM performance, which we attribute to the limited amount of training data in few-shot settings. Accordingly, we consistently set $\ell=10$ for CLIP datasets throughout the paper. For other datasets, updating additional layers leads to better results, so we adopt $\ell=5$. Detailed results are provided in Table 15 in the Appendix.
Adapting Variant Options for PFL. We evaluate the effectiveness of different design choices of MMA in personalized federated learning. In Table 5 (bottom-right), we compare two alternative baselines: treating all adapters as global (Baseline 1) and using the shared adapter as the personalized component while treating the up- and down-projection adapters as global (Baseline 2). As shown, pFedMMA achieves significantly higher local accuracy than both baselines. Moreover, it achieves superior base and novel performance compared to state-of-the-art prompt learning methods, as shown earlier, underscoring its ability to strike a strong balance between personalization and generalization.
4.4 Learning Curves
To further analyze the convergence behavior of pFedMMA, we plot the average local accuracy over communication rounds across five different shot settings in Fig. 4. As shown, pFedMMA consistently attains high accuracy and converges faster than the baselines. These results highlight its ability to effectively balance personalization and generalization, ensuring both rapid and reliable convergence. Detailed results are provided in Fig. 5 in the Appendix.
Figure 4: Accuracy learning curves of pFedMMA and baselines.